# Unsteady aerodynamic theory and experiments of hovering membrane wings

### v1

@@ 4 Concluding remarks
We find very good agreement between the inviscid theory and the experimental results for aeroelastic numbers that are greater than $1.7$ with an angle-of-attack amplitude of $\hat{\alpha}=35^{\circ},55^{\circ}$. Lower values of Ae result with an excessive membrane camber, which leads to vortex formation and flow detachment over the aerofoil. These flow and membrane characteristics violate the assumptions of the theoretical model, and therefore deteriorate the accuracy of the predictive theory.
In general, the good agreement obtained between analytical inviscid theory and measurements of the unsteady lift response of a hovering membrane wing in extreme conditions (high angles of attack) provides a promising validation of theory. In all cases of attached flow and relatively small camber, the unsteady aerodynamic theory provides an accurate prediction of the unsteady lift and can be used to compute the lift coefficient from any experimental or computational data of aerofoil deformation.

### v2

@@ 4 Concluding remarks
We find very good agreement between the inviscid theory and the experimental results for aeroelastic numbers that are greater than $1.7$ with an angle-of-attack amplitude of $\hat{\alpha}=35^{\circ},55^{\circ}$. Lower values of Ae result with an excessive membrane camber, which leads to vortex formation and flow detachment over the aerofoil. These flow and membrane characteristics violate the assumptions of the theoretical model, and therefore deteriorate the accuracy of the predictive theory. Importantly, we find that cases in which the flow remains attached while the maximum camber is substantial still provide good agreement between theory and measurements, indicating the predominant role of the attached flow requirement compared to the small camber condition applied in potential theory.
In general, the good agreement obtained between analytical inviscid theory and measurements of the unsteady lift response of a hovering membrane wing in extreme conditions (high angles of attack) provides a promising validation of theory. In all cases of attached flow and relatively small camber, the unsteady aerodynamic theory provides an accurate prediction of the unsteady lift (mean prediction error between $\Delta\overline{C}_{L_{d}}=0.17$ and $0.27$), and can be used to compute the lift coefficient from any experimental or computational data of aerofoil deformation.